Reports 1-1 of 1 Clear search Modify search
AdV-DET (Commissioning)
gouaty, mwas, swinkels, vardaro - 0:08 Tuesday 08 October 2019 (47140) Print this report
B1 PD2 max current threshold updated

The reason why the DET_MAIN node was stuck in SHUTTER_CHECK_CLOSED (see 47138), was because the node was waiting for the Vbias of the B1_PD2 photodiode to be engaged.But the B1_PD2 photodiode could not be rearmed any more since the last unlock at 06h33 utc this morning.

First we commented in DET_MAIN all the B1_PD2 checks, which allowed to make DET_MAIN happy and to lock the ITF in LN1.

Then we realized that by increasing the maximum current limit of the photodiode from 7.8 to 8.0 in the configuration file of SDB2_dbox_bench_LeftDown, we could rearm the photodiode.

This allowed us to restore Low Noise 3.

The checks involving B1_PD2 in DET_MAIN have been restored.

Comments to this report:
gouaty - 10:51 Tuesday 08 October 2019 (47143) Print this report

This morning during the maintenance, we readjusted the current limit of the B1_PD2 photodiode to the value 7.85, so that it is only slightly above the value that was inducing troubles yesterday. We also tested that the old value (7.8) for the current limit is allowing again to engage the Vbias of the photodiode this morning.

Search Help
×

Warning

×